OpenStudy (anonymous):

find the equation of a line that runs through (7,5) and is parallel to a line with a slope of 3/2

Parth (parthkohli):

\( \color{Black}{\Rightarrow y = {3 \over 2}x + b}\) Find the value of b.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 7 = {3 \over 2}(5) + b}\) \( \color{Black}{\Rightarrow b = 7 - {3 \over 2}{(5)}}\)

OpenStudy (anonymous):

@ParthKohli don't leave with unusual answers try to give fruitful answera

OpenStudy (anonymous):

a. y-7=3/2(x-5 b. y-5=-3/2x-7 c.y-5=3/2(x-7)

Parth (parthkohli):

As I said, find the value of b.

Parth (parthkohli):

And then replace that in the equation I said.

OpenStudy (anonymous):

isn't 7 supposed to be x and 5 is y?

Parth (parthkohli):

\( \color{Black}{\Rightarrow y = mx + b}\) m is 3/2 as it is given. Find the value of b.

Parth (parthkohli):

Oops

OpenStudy (anonymous):

What's the answer?

Parth (parthkohli):

\( \color{Black}{\Rightarrow 5 = {3 \over 2}(7) - b}\)

Parth (parthkohli):

Find the value of b.

Parth (parthkohli):

@Rohangrr did you try to help him?

OpenStudy (anonymous):

@ParthKohli are you helping him in a correct way or just spamming around I have just said you that give a answer in a box. That was a suggestion........ what's the matter of taking that in a very serious tone can u elaborate plzzzz